1/19.5 = 1/x - 1/2x (solve for x)
Answer by stanbon(75887)

1/19.5 = 1/x - 1/2x
Multiply thru by 39x to get:
2x = 39 - 39/2
2x = 39/2
x = 39/4
================
Cheers,
Stan H.